# Which Channel in 6G, Low-rank or Full-rank, more needs RIS?

### v3

@@ I RIS-aided Wireless Network: concept and system
With the development of sixth-generation (6G) communication, there are higher requirements for system capacity, security, and reliability. Reference [1] proposed a secure and precise transmission scheme by combining random subcarrier selection and directional modulation (DM), which enhances system security by concentrating signal power in the small neighbouring region around the legitimate user. Low-cost RISs can be deployed in wireless environments to provide cascaded links, a key technology for improving system performance. In [2], a RIS-assisted DM network was proposed to transmit two confidential data streams from Alice to Bob, achieving twice the secrecy rate of single-stream transmission in the high signal-to-noise ratio region. It is particularly noted that the conventional DM is very suitable for line-of-sight (LoS) channel with rank-one, and only transmits only single bit stream from base station (BS) to user. In [3], a distributed RIS-assisted transmission rate enhancement method in LoS channel was studied to increase channel rank from one to $K$. By dividing a large RIS into multiple small RIS for multi-stream transmission, the transmission rate was improved compared to a large RIS-assisted transmission. In [4], near-field communication (NFC) was discussed as an essential technology for 6G networks, driven by the adoption of extremely large-scale antenna arrays and high frequencies. NFC leverages near-field spherical-wave propagation to enhance beam focusing and improve performance metrics such as signal-to-noise ratio (SNR) and channel capacity. Similarly, reference [5] studied a joint active and passive beamforming design in RIS-aided systems, showing that RIS can achieve comparable performance to massive MIMO systems but with reduced hardware complexity and energy consumption. Furthermore, reference [6] analyzed RIS deployment at cell boundaries to mitigate inter-cell interference, highlighting its role in enhancing spectral efficiency through spatial multiplexing in multicell MIMO systems. In [7], RIS can be combined with relay stations to improve the coverage and transmission rate of the wireless networks. Here, a dual-slot system model with simultaneous deployment of relay stations and RIS was studied, and compared with the single antenna relay network, the performance gain of 85% was achieved in the high SNR region.
@@
As studied above, RIS provides degrees of freedom (DoF) and improves performance compared to the systems without RIS. A joint small-scale spatiotemporal correlation model was proposed and analyzed for DoF under isotropic scattering in [8]. Active RIS can improve the sum of DoFs (sum-DoF) of dual-user MIMO systems with an equal number of antennas at both the transmitting and receiving ends. In [9, 10], the sum-DoF was further studied for any number of receiving and transmitting antennas, and the existence of DoF gain is related to the rate and the number of transmitting and receiving antennas. Based on information theory, the DoF of the joint and independent transmission of information for RIS and transmitter was investigated in[11]. The channel rank of the non-zero direct path can eliminate the phase ambiguity between the RF signal and RIS phase shift caused by multiplicative fading. For the rank-deficient channel, when the rank of the channel from the transmitter to the receiver is large and/or the rank of the channel from the transmitter to the RIS is small, the DoF can be significantly improved [12]. Reference [13] studied the reconfigurable distributed antenna and reflecting surface assisted MIMO system and obtained additional DoF through mode selection of each reflection element. To fully exploit the DoF of RIS, a multi-layer refractive RIS-assisted receiver was proposed [14], and a robust optimization framework of a low-complexity synchronous wireless information and power transmission scheme was designed to enhance receiver performance.
@@ II System Model and DoF Discussion
(a) RoF = 1.
(b) RoF = 2.
(c) RoF = K.
(d) RoF = 0.
(e) RoF = 1.
(f) RoF = K.
Figure 2: Diagram of deploying RIS to increase system degrees of freedom. (a) rank = 1. (b) Increase from rank = 1 to rank = 2. (c) Increase from rank = 1 to rank = $K$. (d) rank = 0. (e) Increase from rank = 0 to rank = 1. (f) Increase from rank = 0 to rank = $K$.
@@ III DoF Improvement and Beamforming in Typical Rank-deficient Channels
Figure 3: The achievable rate versus the number $N$ of RIS elements in scenarios with $\mathbf{H}$.
Figure 4: The achievable rate versus the number $N$ of RIS elements in scenarios where $\mathbf{H}$ is blocked.
To evaluate the rate performance enhancement achieved by DoF increasement due to the introduction of RIS, we will present numerical simulations as follows. System parameters are set as follows: the distances from BS to RIS, RIS to user, and BS to user are 82 m, 28 m, and 100 m, respectively. The maximum ratio transmission (MRT), phase alignment, and maximizing the SNR at transmitter, RIS, and receiver are adopted in our simulation. Fig. 3 plots the achievable rate in the scenarios with $\text{rank}(\mathbf{H})=1$ including RIS-aided and no RIS. Here, the number of RIS elements is 1024 and $K=4$. When $M=64$, the rates of RIS-aided systems in all-LoS and LoS+Raleigh fading channels may about 1.6 times and 2.3 over that of LoS without RIS mainly due to the increase from one to two in system DoF. When the number of antennas at BS is increased from 64 to 128, we attain the same tendency in rate improvement.
@@ IV DoF Boosting and Beamforming via distributed RISs in LoS Channels
Figure 5: Diagram of distributed multi-RIS-aided Wireless Network.
@@
Figure 6: The achievable rate versus the number $N_{j}$ of elements per RIS in scenarios with distributed RISs.
Fig. 6 illustrates the curves of achievable rate versus the number of RIS elements obtained by distributed RISs, where the transmitting beamforming vector, phase adjusting matrix, and receive beamforming vector are designed by null-space-projection, phase alignment, and zero-forcing, respectively. As shown in Fig. 6, as the number of distributed RISs grows from one to four, the achieved rate gains are as follows: 1.4, 2.8, 4.2, and 5.6 times over no RIS under the same power-sum constraint. According to this trend, it can be inferred that increasing the number of RISs will further improve the rate performance.

### v5

# Which Channel in 6G, Low-rank or Full-rank, more needs RIS from a Perspective of DoF?

@@ I RIS-aided Wireless Network: concept and system
With the development of sixth-generation (6G) communication, there are higher requirements for system capacity, security, and reliability. A secure and precise transmission scheme by combining random subcarrier selection and directional modulation (DM) was proposed in [1], which enhances system security by concentrating signal power in the small neighbouring region around the legitimate user. Low-cost RISs can be deployed in wireless environments to provide cascaded links, which is a key technology for improving system performance. It is particularly noted that the conventional DM is very suitable for line-of-sight (LoS) channel with rank-one, which will result in a fact that DM only transmits single bit stream from base station (BS) to user. In [2], a RIS-assisted DM network was proposed to transmit two confidential data streams from Alice to Bob, achieving twice the secrecy rate of single-stream transmission in the high signal-to-noise ratio (SNR) region. In [3], a distributed RISs-assisted transmission rate enhancement method in LoS channel was studied to increase channel rank from one to $K$. By dividing a large RIS into multiple small RISs for multi-stream transmission, the transmission rate was improved compared to a large RIS-assisted transmission. In [4], near-field communication (NFC) was discussed as an essential technology for 6G networks, which leveraged near-field spherical-wave propagation to enhance beam focusing and improve performance metrics such as SNR and channel capacity. Similarly, the authors in [5] studied a joint active and passive beamforming design in RIS-aided systems, which showed that RIS can achieve better performance to massive MIMO systems but with reduced hardware complexity and energy consumption. Furthermore, RIS deployment at cell boundaries was analyzed to mitigate inter-cell interference in [6], which highlighted its role in enhancing spectral efficiency through spatial multiplexing in multicell MIMO systems. In [7], RIS could be combined with relay stations to improve the coverage and transmission rate of the wireless networks. Here, a dual-slot system model with simultaneous deployment of relay stations and RIS was studied, and compared with the single-antenna relay network, a 85% performance gain was achieved in the high SNR region.
@@
As studied above, RIS can improve the performance of wireless network, which is the result of the degrees of freedom (DoF) increased by introducing RIS. A joint small-scale spatiotemporal correlation model was proposed and analyzed for DoF under isotropic scattering in [8]. Active RIS can improve the sum of DoFs (sum-DoF) of dual-user MIMO systems with an equal number of antennas at both the transmitting and receiving ends. In [9, 10], the sum-DoF was further studied for any number of receiving and transmitting antennas, and it was proved that the existence of DoF gain was related to the rate and the number of transmitting and receiving antennas. Based on information theory, the DoF of the joint and independent transmission of information for RIS and transmitter was investigated in[11]. The channel rank of the non-zero direct path can eliminate the phase ambiguity between the radio frequency signal and RIS phase shift caused by multiplicative fading. For the rank-deficient channel, when the rank of the channel from the transmitter to the receiver is large and/or the rank of the channel from the transmitter to the RIS is small, the DoF can be significantly improved [12]. Reference [13] studied the reconfigurable distributed antenna and reflecting surface assisted MIMO system and obtained additional DoF through mode selection of each reflection element. To fully exploit the DoF of RIS, a multi-layer refractive RIS-assisted receiver was proposed [14], and a robust optimization framework of a low-complexity synchronous wireless information and power transmission scheme was designed to enhance receiver performance. In what follows, we will focus on the invesitgation of DoF and rate performance in two typical low-rank channels such as no link and LoS. With the help of RIS, their channel ranks may be augmented from one/zero to full-rank.. This will harvest an extremely large rate gains.
@@ II System Model and DoF Discussion
(a) DoF = 1.
(b) DoF = 2.
(c) DoF = K.
(d) DoF = 0.
(e) DoF = 1.
(f) DoF = K.
Figure 2: Diagram of deploying RIS to increase system DoF: (a) rank = 1; (b) Increase from rank = 1 to rank = 2; (c) Increase from rank = 1 to rank = $K$; (d) rank = 0; (e) Increase from rank = 0 to rank = 1; (f) Increase from rank = 0 to rank = $K$.
@@ III DoF Improvement and Beamforming in Typical Rank-deficient Channels
Figure 3: Curves of achievable rate versus the number $N$ of RIS elements in LoS channel.
Figure 4: Curves of achievable rate versus the number $N$ of RIS elements in scenarios without link.
To evaluate the rate performance enhancement achieved by DoF increasement due to the introduction of RIS, we will present numerical simulations as follows. System parameters are set as follows: the distances from BS to RIS, RIS to user, and BS to user are 82 m, 28 m, and 100 m, respectively. It is not specifically noted that thermal noise powers at RIS and user are -90 dBm and -70 dBm, respectively. The maximum ratio transmission (MRT), phase alignment, and maximizing the SNR at transmitter, RIS, and receiver are adopted in our simulation. Fig. 3 plots the achievable rate in the scenarios with $\text{rank}(\mathbf{H})=1$ including RIS-aided and no RIS. Here, the number of RIS elements is 1024 and $K=4$. When $M=64$, the rates of RIS-aided systems in all-LoS and LoS+Raleigh fading channels may about 1.6 times and 2.3 over that of LoS without RIS mainly due to the increase from one to two in system DoF. When the number of antennas at BS is increased from 64 to 128, we attain the same tendency in rate improvement.
@@ IV DoF Boosting and Beamforming via distributed RISs in LoS Channels
Figure 5: Diagram of distributed multi-RIS-aided Wireless Network in LoS channel.
@@
Figure 6: (a) The achievable rate versus noise power $\sigma_{r}^{2}$ at RIS in scenarios with distributed RISs; (b) The achievable rate versus noise power $\sigma_{u}^{2}$ at user in scenarios with distributed RISs.
To evaluate the rate performance enhancement achieved by DoF increasement due to the introduction of distributed RISs, we will present numerical simulations as follows. System parameters are set as follows: the distances from BS to RIS, RIS to user, and BS to user are 82 m, 28 m, and 100 m, respectively; the number of antennas at BS, the number of RIS elements, and the number of antennas at user are $M=128$, $N=600$, and $K=4$, respectively. Fig. 6a illustrates the curves of achievable rate versus noise power $\sigma_{r}^{2}$ at RIS obtained by distributed RISs with $\sigma_{u}^{2}=-70$ dBm, where the transmitting beamforming vector, phase shifting matrix, and receive beamforming vector are designed by null-space-projection, phase alignment, and zero-forcing, respectively. Observing this figure, we find: as noise power $\sigma_{r}^{2}$ at RIS increases from -120 dBm to -60 dBm, the achieved rate gains over no RIS gradually decrease and converge to zero. When the number of distributed RISs is greater than or equal to 2, i.e., the system DoF increase by 2 or more, the performance is obviously improved compared with no RIS under high noise power, i.e., $\sigma_{r}^{2}=-60$ dBm. As the number of distributed RISs grows from one to four, the achieved rate gains are as follows: about 0.7, 1.9, 2.7, and 3.5 times over no RIS under the same power-sum constraint and $\sigma_{r}^{2}=-90$ dBm. According to this trend, it can be inferred that increasing the number of distributed RISs and reducing noise power $\sigma_{r}^{2}$ at RIS will further improve the rate performance.
Fig. 6b illustrates the curves of achievable rate versus noise power $\sigma_{u}^{2}$ at user obtained by distributed RISs with $\sigma_{r}^{2}=-90$ dBm. From this figure, it is seen: as noise power $\sigma_{u}^{2}$ at user varies from -120 dBm to -60 dBm, the achieved rate gains over no RIS gradually decrease.
In summary, we have the following conclusions: (1) Increasing the number of distributed RISs may boost the value of system rank or DoF such that a significant rate gain can be attained, (2) low-noise RIS and receiver are also crucial for an active RIS-aided wireless network to achieve a high rate gain.
@@ V Open Challenging Problems
  4. 4.
    For a RIS-assisted hybrid beamforming structure, how to jointly design the transmission digital precoding vector, transmission analog weight vector, RIS phase shifting matrix, and reception beamforming vector in order to achieve the maximum DoF of the system?